Outline of Engineering Tasks for UT Jul 9-10-11

Tasks to be done in case SOUL engineering work cannot be accomplished
  1. LUCI-1 Photometric Zeropoint checks - requires clear skies (any seeing)
    • P330E primary
    • 2MASS calibrator fields - multiple visits over different airmasses, z,J,H,K,Ks filters
    • located in /home/observer/Service_Observing/Eng_Jul_2019/PhotCal (RM507 computer)
  2. Poor weather MODS-E task:
    • Examine behavior of open-loop tracking near zenith with MODS.
    • MODS-Blue only photometry (i.e M92) - requires clear skies but any seeing.
  3. LUCI/LBC mixed-mode - 0.5 hr
  4. Compare LBC-B vs LBC-R V-band, r-sloan filters (SX vs DX) - requires clear skies but any seeing
  5. Poor weather Science backup program - any amount of time - targets cover entire night
    • H-band library - K and M stars
    • H-band library - case example of QSO rest-frame 1.6micron CO-bandhead - filler
      • Files located in /home/observer/Service_Observing/Eng_Jul_2019/H-band_Library (RM507 computer)

2019-Jul-09 SX-SOUL-C

SOUL made good use of the 1/2 night. Seeing was stable and <~1" all during the first half, and the team observed a few stars, on-axis, in bins 1 and 2, with LUCI1 through the FeII filter, achieving the predicted Strehl at H. They have detailed notes on their Arcetri wiki, but a brief outline of the night's activities follows:

  1. First, check the re-pointing issue (IT 7640). There is an overshoot in pointing. The sign is ok, but magnitude incorrect.
  2. 1st target: R=7.53 (SOUL measures ~9, delta 1.5). K0 star
  3. 2nd target: R=10.9 (SOUL measures 11.6, delta 0.7). K0 star. Remembered here to offset to sky frame also.
    • It was noted that although both stars are of the same spectral type, there were different deltas between the catalog magnitude and the magnitude measured by the SOUL WFS camera (which has an effective wavelength ~ 8000 angstroms I think).
  4. 3rd target: want 3rd target to be bin 2, which means R=13-14.5 on WFS, which, assuming a ~1.5 mag difference between the catalog magnitude and that measured by the SOUL WFS, means Rcat=11.5-13
    • Tried R=12.9 star (SOUL measures xx, delta xx) AO657 = G125-23 … for some reason, the off-axis star was out of range so we moved on.
    • Tried another R=12.9 star, AO638 = LP449-10. This is an M4 star.(SOUL measures 12.9). Did not observe - too bright for the test.
    • Finally, observed G205-43 (AO681) R=13.5. An M3 star (SOUL measures 13.6).
      • Issue with the PA. NCPA is calibrated so far only for one PA (340 deg). The K-mirror gets stuck at a certain position. It was necessary to unwrap the rotator so as to approach the preset from close to the desired 340 deg, as this would keep the K-mirror off its bad position.
      • Note that for M stars (M3 or M4, at least) the magnitude measured by the SOUL camera matches the R band catalog magnitude.

2019-Jul-10 SX-SOUL-C

SX-SOUL Commissioning continued, this time taking the 2nd half of the night. The theme tonight was vibrations. A 127.4 Hz vibration was seen and that was attributed to the SOUL tip-tilt gain optimization which can excite the resonance frequency of the reference body. That can be resolved, but there was another vibration whose origin was unknown and that is more worrisome. On the LUCI1 images, the axis of elongation and amplitude changed with LFBG rotator angle.

  1. 1st target: AO722 = HD 200371 K5 R=7.32 (SOUL 8.7)
    • set subwin to 128 128 (that’s the minimum and gives 0.15672 s exptime) … still 53000 counts
    • angle was at 103. Need to set PA to get rotator angle = 340.
    • The modal gain optimization seems to excite a vibration at 127.4 Hz. This is non-linear and therefore seems to be mechanical rather than due to hitting a resonance. It could be that the high frequency of SOUL excites a resonance in the reference body.
    • Changing ROTANGLE - going from 103 down to smaller values, -13 deg Bad for K mirror. Need to move LUCI1 rotator to 300 and go from there.
    • But now star is on a bad pixel! Need to make an offset. The loop was paused, but optimize gain did not stop/start so it accumulated some things. The star looks like a “+” now.
  2. 2nd target: AO748 = G212=16 M4 R=14.4(SOUL 14.0) H=10.006 (54000 counts - full frame with 2.51-sec exptime)
    • Optimize gain failed. Preset had to be cancelled and resent.
    • Try H band. Full frame 947, 1098 ok
    • 09:36 - very bad active optics WFS image - had offloaded a lot of aberr onto M1. After runAO failed. Steve had turned off sending Zernikes. Now Steve clearing M1 active optics only.
    • Now we are vignetting on M3 - the guide star is near the edge of the field. ProbeX, ProbeY = 133.94, 122.15 [mm,mm]. But good enough - AO will take over.
    • But actually, this star is too bright. SOUL estimates magnitude 14.0 and we're still in bin2.
  3. 3rd target: AO715 = M5 LP 458-43 star R=15.4 (SOUL 15.7) H=11.247
    • 10:09 — AO software went down, AO-SX AOSL not connected to supervisor. adam supervisor and fast diagnostics went down. No one did an adsec stop. Unclear what caused it to go down.
    • As a consequence of AO SW glitch, AO offloaded a lot of junk to Active Optics. Steve clearing ActOpt. Doug & SOUL team restarting wunit processes.
    • 10:16 ready to go again.
    • H-band. subwin 256x256.
    • Lots of vibration!! About a 10:1 axis ratio. This is an input vibration - it is independent from the gain, so not self-inflicted. 
Nothing on OVMS that is different from earlier.
    • Currently at rotangle = 316. Doug suggests changing rotangle - he’s seen vibration earlier that is rotator angle dependent. Elongation ~45 deg.
    • Next tried rotangle~332 degrees, it’s better, but there’s still some vibration. It appears to be along the same axis as before. (but maybe shallower? ~22 deg).
    • We’ll rotate another 20 deg. Angle seems to have changed to about ~22 now.
    • 11:18. 360 deg. Axis of vibration is now horizontal.
    • Greg is offsetting by 10” for sky. Failed - fast link comm problem. Success.
    • 11:33 Taking exposures. Bin 3: 54 modes.
    • 11:41: Twilight is here and sky background is increasing. Need to go to a brighter star.
  4. 4th target: AO795 = BD+19 4896 K2 H=5.777 G=8.29 (G=Gaia broad band mag)
    • rotangle=360
    • OVMS shows 127.4 Hz vibration - SOUL induces this vibration and OVMS picks it up.
    • Is this a double star? It looks that way on LUCI1 images. Simbad shows it as K2 D~ (is the D~ double - not obvious it is catalogued as a double).

11:58 UT: Shutting down. Warming up OCAM

UT Jul 11 2019

UT 06:43 - closing up dome - thick clouds overhead - handover is 00:25 local time

UT 06:59 - configuring for LUCI-1 SX - iLocater officially handing over early - LUCI software started, 2x test darks taken (luci1.20190711.0001-0002) - init all ran

UT 07:26 - cal unit deployed, lights out, closed dome testing proceeding.

UT 08:07 - noticing some vibrations on the telescope affecting the PSF

-collecting data starting with luci1.20190711.0006.fits

-submitted IT 7780 - motion in LUCI-1 with AO star image

UT 09:34 - still closed - testin tip-tilt offsets

UT 10:12 - finished with offset test - moving ARGOS cal unit out of the way. Preparing to open - allsky shows clear(ish) patches overhead

--first target AO 795 R=7.8 K2 (H-mag ~ 6)


subframe set up 256x256, FeII filter first frame luci1.20190711.0063


vibrations again, PSF looks elongated

UT 10:51 - PSF quality is terrible, images become like cigars

UT 11:19 - trying R~14mag - which with conditions should be R~16 M3.5 star

--having difficulty, could be at -12 twilight the cause of the problem

-took a couple of test images

-Putting LUCI-1 into safe config, done for the night - closing up

-- BarryRothberg - 05 Jul 2019
Topic revision: r8 - 29 Sep 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback